stranger2:

Geometry @jhonny9 @snowflake0531

stranger2:

snowflake0531:

hehe, this is an isosceles triangle

snowflake0531:

so x and y are equal lol, i think

stranger2:

ohh

snowflake0531:

and i forgot to explain why they're isosceles

snowflake0531:

it's since sin and cos are both something over hypotenus, and since the something is equal, the two angles are equal, so it's isosceles, if that makes sense lol

jhonyy9:

sin x = cos y op/hyp = adj/hyp op = adj 

jhonyy9:

given m<x = 72° m<y= ?

stranger2:

wait where is m

jhonyy9:

measure of angle

jhonyy9:

measure of angle x noted by m<x =

jhonyy9:

ok ?

stranger2:

ok-

jhonyy9:

can you find the y measure ?

stranger2:

no idk

jhonyy9:

look pls on my posted image

jhonyy9:

the sum of angles measures of a triangle equal how many ?

jhonyy9:

equal 180 degrees yes ?

stranger2:

yes

jhonyy9:

on my posted image there is a right angle with how many degrees ?

jhonyy9:

a right angle has 90 degrees

jhonyy9:

so you know that there is an angle with measure of 72 this is angle x and the right angle with measure of 90 degrees and the total need be 180 how you calcule the 3rd angle measure ?

jhonyy9:

y +72 +90 = 180 find y

jhonyy9:

can you calcule the y value from this equation ?

jhonyy9:

y +162 =180 y = 180 - 162 y = ?

jhonyy9:

ok ?

stranger2:

so- 18

jhonyy9:

yes and what choice maybe correct ?

stranger2:

A?

jhonyy9:

perfect

stranger2:

thankssss

jhonyy9:

understood it ?

stranger2:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 understood it ? \(\color{#0cbb34}{\text{End of Quote}}\) yes ty so much

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @stranger2 thankssss \(\color{#0cbb34}{\text{End of Quote}}\) anytime

Downpour17:

Noice Jhonyy, you is big brain